ANALYTICAL EXPRESSIONS OF FREQUENCIES, DIPOLE STRENGTHS AND ROTATIONAL STRENGTHS FOR OVERTONE TRANSITIONS IN A DISSYMMETRIC HCCH FRAGMENT

\underline{ROBERTO GANGEMI}, G. LONGHI, F. LEBON, S. ABBATE, Dipartimento di Scienze Biomediche e Biotecnologie, Universitá di Brescia, Viale Europa 11, 25123 Brescia, Italy.

In a recent paper ^{a} we derived useful analytical expressions for frequencies, dipole strengths and rotational strengths for CH-stretching transitions Δv = 1, 2, 3 in a two degrees of freedom model of an HCCH fragment endowed with C2symmetry. The model includes mechanical and electrical anharmonicities ^{b}, which are dealt with by second order Van Vleck perturbation theory ^{c}; expressions are derived through a computer algebraic manipulator. This dynamical study has been pursued quite efficiently also by other authors ^{d} ^{e}, with Hamiltonians written in second quantization formalism. Herein we report on an update of our approach, that includes ideas derived from refs. (d) and (e), and allows us to deal with CH-stretching overtones dipole and rotational strengths up to Δv = 4, for a more general fragment deprived of C2-symmetry. In the latter case a 1:1 resonance is noticed in addition to the 2:2 Darling-Dennison resonance.

aS. Abbate, R. Gangemi, G. Longhi, J. Chem. Phys. \underline{117}, 7575(2002)

bG. Longhi, L. Ricard, S. Abbate, G. Zerbi, J. Chem. Phys. \underline{88}, 6733(1988)

cG. Amat, H. H. Nielsen, G. Tarrago, Rotation-Vibrations spectra of Polyatomic Molecules, M. Dekker, New York

dL. Halonen, Adv. Chem. Phys. \underline{104}, 41(1998)

eX. G. Wang, E. L. Sibert III, J. Chem. Phys. \underline{113}, 5384(2000)